# Thompson sampling for zero-inflated count outcomes with an application to the Drink Less mobile health study

### v1

@@ 4 The TS-Count Strategy
2       Update the posterior distribution $Q_{t}(\gamma)$ and $Q_{t}(\beta)$ using data $D_{t-1}$ Observe context $X_{t}$ Sample $\tilde{\gamma}_{t}$ from the updated posterior distribution $Q_{t}(\gamma)$ Sample $\tilde{\beta}_{t}$ from the updated posterior distribution $Q_{t}(\beta)$ Compute $A_{t}=\arg\max_{a\in\mathcal{A}}\left[1-\operatorname{sigmoid}(\phi(a,X_{t})^{\top}\tilde{\gamma}_{t})\right]\exp\left\{\phi(a,X_{t})^{\top}\tilde{\beta}_{t}\right\}$ Offer intervention $A_{t}$ and observe the associated outcome $Y_{t}$
@@
This algorithm operates through a sequential five-step process at each decision time $t$. First, a prior distribution $Q_{0}(\beta)$ is assumed for the underlying model parameter $\beta^{*}$. Subsequently, the algorithm proceeds to update the posterior distribution $Q_{t}(\beta)$ by integrating the data collected before time $t$, say $D_{t-1}$. Following this, the algorithm derives a Bayesian estimate $\tilde{\beta}_{t}$ for $\beta^{*}$ by sampling from the posterior distribution $Q_{t}(\beta)$. This process is grounded in the Bayesian framework, which treats the true parameter $\beta^{*}$ as a random variable; thus, at time $t$, both the Bayesian estimate $\tilde{\beta}_{t}$ and the model parameter $\beta^{*}$ are characterized by a posterior distribution $Q_{t}(\beta)$. The subsequent phase involves selecting an intervention $A_{t}$ that maximizes the estimated expected proximal outcome under the Bayesian estimate $\tilde{\beta}_{t}$. Finally, $A_{t}$ is offered to the user, leading to the observation of a proximal outcome $Y_{t}$. It is important to note that for TS-ZIP and TS-ZINB, an additional step is included wherein $\tilde{\gamma}_{t}$ is sampled from its respective posterior distribution to aid in the intervention selection process, thereby introducing an extra layer of randomness into the algorithm.
The proposed algorithm, which updates the posterior distribution $Q_{t}(\gamma)$ and $Q_{t}(\beta)$ through a Bayesian framework, benefits significantly from prior assumptions about the underlying model parameter $\gamma^{*}$ and $\beta^{*}$. While a normal prior is frequently utilized in the estimation of regression coefficients due to its desirable properties (van Zwet, 2019), its application within count data models presents computational challenges. Specifically, the normal prior is non-conjugate for such models, thereby complicating the derivation of closed-form posterior distributions. This inherent complexity necessitates the adoption of advanced computational methods. Markov chain Monte Carlo (MCMC) techniques become indispensable in such scenarios to approximate the posterior distribution effectively. However, MCMC can be computationally demanding and rather time-consuming, particularly in online settings where the algorithm updates at each time $t$ (Russo et al., 2018). As such, we turn to more efficient methods for approximate sampling from posterior distributions.
@@ 4.1 Laplace Approximation for Tractability
In this section, we address the computational complexity associated with the posterior distributions in the TS-Count framework. Our approach involves adopting the Laplace approximation, a technique that simplifies the posterior distribution by approximating it with a normal distribution (Russo et al., 2018). This approximation allows us to generate samples from a normal distribution, which serves as proxies for samples from the actual posterior distribution. As Abeille and Lazaric (2017) have indicated, when a distribution possesses appropriate concentration and anti-concentration characteristics, it ensures minimal regret in the approximation. Specifically, the stochastic estimate $\tilde{\beta}_{t}$ is designed to strike a balance between being sufficiently far from the MLE $\hat{\beta}_{t}$ (anti-concentration) to ensure enough exploration, and being adequately concentrated around $\hat{\beta}_{t}$ (concentration) to ensure adequate exploitation.
At each decision time $t$, a Bayesian estimate $\tilde{\beta}_{t}$ is sampled from a Gaussian distribution $\mathcal{N}(\hat{\beta}_{t},\alpha_{\beta}^{2}\mathcal{I}_{t}^{-1}(\hat{\beta}_{t}))$. Here, $\hat{\beta}_{t}$ denotes the MLE, and $\mathcal{I}_{t}(\hat{\beta}_{t})$ represents the Fisher information, with $\alpha_{\beta}$ serving as a tuning parameter. Similarly, for our implementations of TS-ZIP and TS-ZINB, we sample $\tilde{\gamma}_{t}$ from a Gaussian distribution $\mathcal{N}(\hat{\gamma}_{t},\alpha_{\gamma}^{2}\mathcal{I}_{t}^{-1}(\hat{\gamma}_{t}))$, where $\alpha_{\gamma}$ is also a tuning parameter.The pseudocode for applying TS-Poisson or TS-NB with Laplace approximation is detailed in Algorithm 3. In contrast, Algorithm 4 presents the pseudocode of TS-ZIP or TS-ZINB with Laplace approximation. The use of Laplace approximation in these algorithms aids in efficiently approximating posterior distributions. The tuning parameters, $\alpha_{\beta}$ and $\alpha_{\gamma}$, are critical in controlling the exploration-exploitation tradeoff and must be chosen carefully.
@@
1 Initialization: Randomly choose $A_{t}\in\mathcal{A}$ for $t\leq\tau$ and collect the observations $D_{\tau}$ for $t=\tau+1,\cdots,T$ do
       Compute the MLE $\hat{\beta}_{t}$ by maximizing the likelihood function /* TS-Poisson: Equation (1) */
@@
2       Get $\mathcal{I}_{t}(\hat{\beta}_{t})$ as a byproduct of step 3 Sample $\tilde{\beta}_{t}$ from $\mathcal{N}(\hat{\beta}_{t},\alpha_{\beta}^{2}\mathcal{I}_{t}^{-1}(\hat{\beta}_{t}))$ Observe context $X_{t}$ and compute $A_{t}=\arg\max_{a\in\mathcal{A}}\exp\left\{\phi(a,X_{t})^{\top}\tilde{\beta}_{t}\right\}$ Offer intervention $A_{t}$ and observe the associated outcome $Y_{t}$
@@
1 Initialization: Randomly choose $A_{t}\in\mathcal{A}$ for $t\leq\tau$ and collect the observations $D_{\tau}$ for $t=\tau+1,\cdots,T$ do
       Compute the MLEs $\hat{\gamma}_{t}$ and $\hat{\beta}_{t}$ by maximizing the likelihood function /* TS-ZIP: Equation (3) */
@@
2       Get $\mathcal{I}_{t}(\hat{\gamma}_{t})$ and $\mathcal{I}_{t}(\hat{\beta}_{t})$ as a byproduct of step 3 Sample $\tilde{\gamma}_{t}$ from $\mathcal{N}(\hat{\gamma}_{t},\alpha_{\beta}^{2}\mathcal{I}^{-1}_{t}(\hat{\gamma}_{t}))$ Sample $\tilde{\beta}_{t}$ from $\mathcal{N}(\hat{\beta}_{t},\alpha_{\gamma}^{2}\mathcal{I}^{-1}_{t}(\hat{\beta}_{t}))$ Observe context $X_{t}$ and compute $A_{t}=\arg\max_{a\in\mathcal{A}}\left[1-\operatorname{sigmoid}(\phi(a,X_{t})^{\top}\tilde{\gamma}_{t})\right]\exp\left\{\phi(a,X_{t})^{\top}\tilde{\beta}_{t}\right\}$ Offer intervention $A_{t}$ and observe the associated outcome $Y_{t}$
@@ 4.2 Performance Metrics
where $h(\phi(a,X_{t})^{\top};\eta^{*})$ denotes the expected proximal outcome for intervention $a$ at time $t$, $\eta^{*}$ denotes the vector of true model parameters. For TS-Poisson and TS-NB, $h(\phi(a,X_{t})^{\top};\eta^{*})=\exp\left\{\phi(a,X_{t})^{\top}\beta^{*}\right\}$, while for TS-ZIP and TS-ZINB, $h(\phi(a,X_{t})^{\top};\eta^{*})=\left[1-\operatorname{sigmoid}(\phi(a,X_{t})^{\top}\gamma^{*})\right]\allowbreak\exp\left\{\phi(a,X_{t})^{\top}\beta^{*}\right\}$. Note that, the outer expectation is taken over random variables $A_{t}$ and $X_{t}$, which may depend on the randomness in rewards and/or in the algorithm.
@@
The concept of regret is central to measuring the performance of a contextual bandit algorithm. It quantifies the expected difference between the expected proximal outcome generated by the optimal intervention and the intervention selected by the learning algorithm. A lower regret value indicates a better performance of the learning algorithm, with zero regrets being achieved by an oracle algorithm that always selects the optimal intervention. By focusing on minimizing regret, we aim to develop an approach that can improve user engagement with the Drink Less app.
5 Theoretical Results
In this section, we present the theoretical guarantees, i.e., regret bounds, for the proposed algorithms. Previous studies, such as Russo and Van Roy (2014), have shown that regret analysis for generalized linear bandits can be extended to TS with a Poisson model (TS-Poisson). However, such an extension is not straightforward for TS-NB, TS-ZIP, and TS-ZINB. The complexity arises primarily due to the undetermined overdispersion parameter in NB regression and the parameters governing the zero components in ZIP and ZINB.
To facilitate our analysis, we introduce some new notations. The norm of a vector $x$ is denoted as $|x|=\sqrt{x^{\top}x}$, and $\tilde{O}$ denotes the big-O notation up to logarithmic factors in time $T$. We define $G_{t}=\sum_{i=1}^{t-1}\phi(A_{i},X_{i})\phi(A_{i},X_{i})^{\top}$ as the unweighted Hessian matrix at decision time $t$. For any positive semi-definite matrix $M$, $\lambda_{\min}(M)\geq 0$ denotes the minimum eigenvalue of $M$. Additionally, $\Delta_{\max}$ represents the maximum gap, across all actions $a$ in the action set $\mathcal{A}$, between the optimal expected proximal outcome and that of a selected intervention $a$, namely, $\Delta_{\max}=\max_{a\in\mathcal{A}}\Delta_{a}$. For establishing the regret bounds of TS-Count variants, our analysis depends on several key assumptions, which will be detailed subsequently.
@@ 5 Theoretical Results
For all time $t$ and any intervention $a$, we have $\left\|\phi(a,X_{t})\right\|\leq 1$. In addition, we have $\|\gamma^{*}\|\leq 1$ and $\|\beta^{*}\|\leq 1$.
@@
Assumption 1 is used to make the regret bounds scale-free for convenience; otherwise, we may assume $\left\|\phi(a,X_{t})\right\|\leq C$, $\|\gamma^{*}\|\leq C$, and $\|\beta^{*}\|\leq C$ instead, and the regret bound would increase by a factor of $C$. In practice, this assumption can also be satisfied by scaling the data. Assumption 2 restricts the magnitude of the proximal outcome, which is adapted from the generalized linear bandit literature (Filippi et al., 2010). We make an additional Assumption 3 to facilitate the regret analysis for TS-NB and TS-ZINB.
Theorem 1.
Suppose we run TS-NB for a total of $T$ time points. Under Assumptions 1-3 (also satisfying certain mild regularity conditions defined in the Supplement), the Bayesian regret of the algorithm is upper-bounded as
[…]
@@
Theorem 1 establishes $\tilde{O}(d\sqrt{T})$ Bayesian regret, which coincides with the Bayesian regret bound for generalized linear TS (Russo and Van Roy, 2014). A detailed proof of Theorem 1 is deferred to Section A.1 of the Supplement.
Theorem 2.
Suppose we run TS-ZIP for a total of $T$ time points. Under Assumptions 1-3 (also satisfying certain mild regularity conditions defined in the Supplement), the Bayesian regret of the TS-ZIP algorithm is upper-bounded as
[…]
$+$ $\left[2qY_{\max}\kappa_{1,\min}^{-1}\sqrt{d\log(T/d)+4\log T}\right]\sqrt{2dT\log\left(\frac{T}{d}\right)}$
$+$ $\left[2q\kappa_{2,\min}^{-1}\sqrt{d\log(T/d)+4\log T}\right]\sqrt{2dT\log\left(\frac{T}{d}\right)},$
@@
Theorem 2 also establishes $\tilde{O}(d\sqrt{T})$ Bayesian regret for TS-ZIP. The proof is deferred to Sections A.2 of the Supplement. The Bayesian regret for TS-ZINB has the same form as Equation (6), but under slightly different conditions, and the proof can be found in Section A.3 of the Supplement.
In the second part of our analysis, we turn our focus to the evaluation of frequentist regret associated with the proposed algorithms. Frequentist regret, in contrast to Bayesian regret, provides a measure of performance under the assumption that true parameters of the model are fixed but unknown, rather than being treated as random variables (Hamidi and Bayati, 2020). Frequentist regret is particularly significant as it offers a robust metric that emphasizes the worst-case or minimax regret. This approach involves bounding the maximum regret possible for any given instance of the problem, thereby providing insights into the algorithm’s performance in the most challenging conditions (Agrawal and Jia, 2017).
Theorem 3.
@@
$\mathcal{R}(T,\beta^{*})\leq s(c_{1}+c_{2})\left(1+\frac{2}{0.15-1/T}\right)\times\sqrt{2dT\log(T/d)}+(\tau+2)\Delta_{\max},$
where
$c_{1}$ $=$ $Y_{\max}\kappa_{\min}^{-1}\sqrt{d\log(T/d)+2\log T},$
$c_{2}$ $=$ $c_{1}\sqrt{2\kappa_{\min}^{-1}\kappa_{\max}\log(KT)}.$
Theorem 3 establishes $\tilde{O}(d\sqrt{T\log K})$ regret, which coincides with the regret bound for TS in generalized linear bandits (Kveton et al., 2020). We defer the proof of Theorem 3 to Section B.1 of the Supplement.
Theorem 4.

### v2

@@ 6      Offer intervention $A_{t}$ and observe the associated outcome $Y_{t}$
2       Update the posterior distribution $Q_{t}(\gamma)$ and $Q_{t}(\beta)$ using data $D_{t}$
3      Observe context $X_{t}$
4      Sample $\tilde{\gamma}_{t}$ from the updated posterior distribution $Q_{t}(\gamma)$
5      Sample $\tilde{\beta}_{t}$ from the updated posterior distribution $Q_{t}(\beta)$
6      Compute $A_{t}=\arg\max_{a\in\mathcal{A}}\left[1-\operatorname{sigmoid}(\phi(a,X_{t})^{\top}\tilde{\gamma}_{t})\right]\exp\left\{\phi(a,X_{t})^{\top}\tilde{\beta}_{t}\right\}$
7      Offer intervention $A_{t}$ and observe the associated outcome $Y_{t}$
8 end for
@@ 7      Offer intervention $A_{t}$ and observe the associated outcome $Y_{t}$
These algorithms operate through a sequential five-step process at each decision time $t$. First, a prior distribution $Q_{0}(\beta)$ is assumed for the underlying model parameter $\beta^{*}$. Subsequently, the algorithm proceeds to update the posterior distribution $Q_{t}(\beta)$ by integrating the data collected up to time $t$, say $D_{t}$. Following this, the algorithm derives a Bayesian estimate $\tilde{\beta}_{t}$ for $\beta^{*}$ by sampling from the posterior distribution $Q_{t}(\beta)$. This process is grounded in the Bayesian framework, which treats the true parameter $\beta^{*}$ as a random variable; thus, at time $t$, both the Bayesian estimate $\tilde{\beta}_{t}$ and the model parameter $\beta^{*}$ are characterized by a posterior distribution $Q_{t}(\beta)$. The subsequent phase involves selecting an intervention $A_{t}$ that maximizes the estimated expected proximal outcome under the Bayesian estimate $\tilde{\beta}_{t}$. Finally, $A_{t}$ is offered to the user, leading to the observation of a proximal outcome $Y_{t}$. It is important to note that for both TS-ZIP and TS-ZINB, an additional step is included wherein $\tilde{\gamma}_{t}$ is sampled from its respective posterior distribution to aid in the intervention selection process, thereby introducing an extra layer of randomness into the algorithm.
The proposed algorithms, which update the posterior distribution $Q_{t}(\gamma)$ and $Q_{t}(\beta)$ through a Bayesian framework, benefit significantly from prior assumptions about the underlying model parameter $\gamma^{*}$ and $\beta^{*}$. While a normal prior is frequently utilized in the estimation of regression coefficients due to its desirable properties (van Zwet, 2019), its application within count data models presents computational challenges. Specifically, the normal prior is non-conjugate for such models, thereby complicating the derivation of closed-form posterior distributions. This inherent complexity necessitates the adoption of advanced computational methods. Markov chain Monte Carlo (MCMC) techniques become indispensable in such scenarios to approximate the posterior distribution effectively. However, MCMC can be computationally demanding and rather time-consuming, particularly in online settings where the algorithm updates very frequently (Russo et al., 2018). As such, we turn to more efficient methods for approximate sampling from posterior distributions.
@@ 4.1 Laplace Approximation for Tractability
In this section, we address the computational complexity associated with the posterior distributions in the TS-Count framework. Our approach involves adopting the Laplace approximation, a technique that simplifies the posterior distribution by approximating it with a normal distribution (Russo et al., 2018). As Abeille and Lazaric (2017) have indicated, when a distribution possesses appropriate concentration and anti-concentration characteristics, it ensures minimal regret in the approximation. Specifically, the stochastic estimate $\tilde{\beta}_{t}$ is designed to strike a balance between being sufficiently far from the MLE $\hat{\beta}_{t}$ (anti-concentration) to ensure enough exploration, and being adequately concentrated around $\hat{\beta}_{t}$ (concentration) to ensure adequate exploitation.
At each decision time $t$, a Bayesian estimate $\tilde{\beta}_{t}$ is sampled from a Gaussian distribution $\mathcal{N}(\hat{\beta}_{t},\alpha_{\beta}^{2}\mathcal{I}_{t}^{-1}(\hat{\beta}_{t}))$. Here, $\hat{\beta}_{t}$ denotes the MLE, and $\mathcal{I}_{t}(\hat{\beta}_{t})$ represents the Fisher information, with $\alpha_{\beta}$ serving as a tuning parameter. Similarly, for our implementations of TS-ZIP and TS-ZINB, we additionally sample $\tilde{\gamma}_{t}$ from a Gaussian distribution $\mathcal{N}(\hat{\gamma}_{t},\alpha_{\gamma}^{2}\mathcal{I}_{t}^{-1}(\hat{\gamma}_{t}))$, where $\alpha_{\gamma}$ is also a tuning parameter. The pseudocode for applying TS-Poisson or TS-NB with Laplace approximation is detailed in Algorithm 3. In contrast, Algorithm 4 presents the pseudocode of TS-ZIP or TS-ZINB with Laplace approximation. The implementation of Laplace approximation in these algorithms facilitates efficient sampling of the stochastic estimates $\tilde{\beta}_{t}$ and $\tilde{\gamma}_{t}$. The tuning parameters, $\alpha_{\beta}$ and $\alpha_{\gamma}$, are critical in controlling the exploration-exploitation tradeoff and must be chosen carefully (see the proofs of Theorems 5.3 - 5.4 in the Supplement for more details).
@@
1 Initialization: Randomly choose $A_{t}\in\mathcal{A}$ for $t\leq\tau$ and collect the observations $D_{\tau+1}$
2for $t=\tau+1,\cdots,T$ do
3
4      Compute the MLE $\hat{\beta}_{t}$ by maximizing the log-likelihood function using data $D_{t}$
      /* TS-Poisson: Equation (1) */
@@ 4      Compute the MLE $\hat{\beta}_{t}$ by maximizing the log-likelihood function using data $D_{t}$
5
6      Get $\mathcal{I}_{t}(\hat{\beta}_{t})$ as a byproduct of step 3
7      Sample $\tilde{\beta}_{t}$ from $\mathcal{N}(\hat{\beta}_{t},\alpha_{\beta}^{2}\mathcal{I}_{t}^{-1}(\hat{\beta}_{t}))$
8      Observe context $X_{t}$ and compute $A_{t}=\arg\max_{a\in\mathcal{A}}\exp\left\{\phi(a,X_{t})^{\top}\tilde{\beta}_{t}\right\}$
9      Offer intervention $A_{t}$ and observe the associated outcome $Y_{t}$
10 end for
@@ 9      Offer intervention $A_{t}$ and observe the associated outcome $Y_{t}$
1 Initialization: Randomly choose $A_{t}\in\mathcal{A}$ for $t\leq\tau$ and collect the observations $D_{\tau+1}$
2for $t=\tau+1,\cdots,T$ do
3
4      Compute the MLEs $\hat{\gamma}_{t}$ and $\hat{\beta}_{t}$ by maximizing the log-likelihood function using data $D_{t}$
      /* TS-ZIP: Equation (3) */
@@
5
6      Get $\mathcal{I}_{t}(\hat{\gamma}_{t})$ and $\mathcal{I}_{t}(\hat{\beta}_{t})$ as a byproduct of step 3
7      Sample $\tilde{\gamma}_{t}$ from $\mathcal{N}(\hat{\gamma}_{t},\alpha_{\beta}^{2}\mathcal{I}^{-1}_{t}(\hat{\gamma}_{t}))$
8      Sample $\tilde{\beta}_{t}$ from $\mathcal{N}(\hat{\beta}_{t},\alpha_{\gamma}^{2}\mathcal{I}^{-1}_{t}(\hat{\beta}_{t}))$
9      Observe context $X_{t}$ and compute $A_{t}=\arg\max_{a\in\mathcal{A}}\left[1-\operatorname{sigmoid}(\phi(a,X_{t})^{\top}\tilde{\gamma}_{t})\right]\exp\left\{\phi(a,X_{t})^{\top}\tilde{\beta}_{t}\right\}$
10      Offer intervention $A_{t}$ and observe the associated outcome $Y_{t}$
11 end for
@@ 10      Offer intervention $A_{t}$ and observe the associated outcome $Y_{t}$
We define $G_{t}=\sum_{i=1}^{t-1}\phi(A_{i},X_{i})\phi(A_{i},X_{i})^{\top}$ as the unweighted Hessian matrix at decision time $t$. TS-Count with normal approximation starts with an initial exploration time period $\tau$, which consists of delivering interventions $A_{1},A_{2},\cdots,A_{\tau}$ with uniform randomization probabilities. This can ensure the invertibility of $G_{t}$ for all $t\geq\tau$. The necessity is motivated by the classical likelihood theory: as the sample size $t\rightarrow\infty$, the MLE $\hat{\beta}_{t}$ is consistent and asymptotically normal, i.e., $\hat{\beta}_{t}-\beta^{*}\rightarrow\mathcal{N}(0,\mathcal{I}_{t}^{-1}(\beta^{*}))$. The invertibility of $G_{t}$ can ensure that $\mathcal{I}_{t}$ is invertible and a unique MLE is attainable. Proposition 4.1 provides a way for theoretically specifying the length of the initial exploration time.
Proposition 4.1.
This is an adaptation of Proposition 1 in Li, Lu and Zhou (2017) and Oh and Iyengar (2019) to TS-Count, and the proof can be found in the same sources. Note that the i.i.d. assumption on $X_{t}$ is only required during the initial exploration period to ensure $G_{t}$ is invertible. Subsequently, $X_{t}$ can be chosen adversarially, without assuming any structure on it, as long as $||X_{t}||$ is bounded.
Remark 4.1.
In terms of the Drink Less app, Proposition 4.1 provides guidance on the number of initial observations required to ensure the validity of the learning algorithm for intervention delivery. Typically, we select $\tau$ such that the minimum eigenvalue of $G_{\tau}$ satisfy $\lambda_{\min}(G_{\tau})\geq 1$, aligning with setting $B=1$ in Proposition 4.1. Additionally, when designing a real mobile health study, it is important to ensure that the length of the study is longer than the required initial exploration length. This consideration needs to be taken into account before starting the study to ensure sufficient data collection beyond the exploration phase.
An alternative approach involves the use of regularization techniques. Specifically, we suggest modifying Line 3 in Algorithms 3 and 4 to include a ridge penalty in the objective function, $\lambda\|\beta\|_{2}^{2}$, mirroring the effect of a normal prior within a Bayesian context. This is of particular interest during the implementation of algorithms for the Drink Less case study, as discussed in Section 7.
@@ 4.2 Performance Metrics
where $h(\phi(a,X_{t})^{\top};\eta^{*})$ denotes the expected proximal outcome for intervention $a$ at time $t$, $\eta^{*}$ denotes the vector of the true model parameters. For TS-Poisson and TS-NB, $h(\phi(a,X_{t})^{\top};\eta^{*})=\exp\left\{\phi(a,X_{t})^{\top}\beta^{*}\right\}$, while for TS-ZIP and TS-ZINB, $h(\phi(a,X_{t})^{\top};\eta^{*})=\left[1-\operatorname{sigmoid}(\phi(a,X_{t})^{\top}\gamma^{*})\right]\allowbreak\exp\left\{\phi(a,X_{t})^{\top}\beta^{*}\right\}$. Note that, the outer expectation is taken over random variables $A_{t}$ and $X_{t}$, which may depend on the randomness in rewards and in the algorithm.
@@
The concept of regret is central to measuring the performance of a contextual bandit algorithm. It quantifies the expected difference between the expected proximal outcome generated by the optimal intervention and the intervention selected by the learning algorithm. A lower regret value indicates a better performance of the learning algorithm, with zero regrets being achieved by an oracle algorithm that always selects the optimal intervention.
Bayesian regret is less informative than its frequentist counterpart because a bound on $\mathcal{R}_{Bayes}(T)$ does not necessarily translate to a meaningful bound on $\mathcal{R}(T,\eta^{*})$. However, the converse is true. Frequentist regret, by focusing on the worst-case or minimax regret, offers a robust measure that bounds the maximum possible regret for any specific instance of the problem, thus providing valuable insights into the algorithm’s performance under the most challenging conditions (Agrawal and Jia, 2017).
5 Regret Analysis
In this section, we present the theoretical guarantees, i.e., regret bounds, for the proposed algorithms. Previous studies, such as Russo and Van Roy (2014), have shown that regret analysis for generalized linear bandits can be extended to TS with a Poisson model (TS-Poisson). However, such an extension is not straightforward for TS-NB, TS-ZIP, and TS-ZINB. The complexity arises primarily due to the undetermined overdispersion parameter in NB regression and the parameters governing the zero components in ZIP and ZINB. In what follows, we first establish Bayesian regret bounds for Algorithms 1 and 2, assuming the posterior distributions $Q_{t}(\beta)$ and $Q_{t}(\gamma)$ are known. We then transition to examining Algorithms 3 and 4, where we derive more robust frequentist regret bounds. It is important to note that the concentration results obtained from the Bayesian regret analysis are crucial for formulating the frequentist regret bounds in Algorithms 3 and 4.
To facilitate our analysis, we introduce some new notations. The norm of a vector $x$ is denoted as $|x|=\sqrt{x^{\top}x}$, and $\tilde{O}$ denotes the big-$O$ notation up to logarithmic factors in time $T$. For any positive semi-definite matrix $M$, $\lambda_{\min}(M)\geq 0$ denotes the minimum eigenvalue of $M$. Additionally, $\Delta_{\max}$ represents the maximum gap across all actions $a$, namely, $\Delta_{\max}=\max_{a\in\mathcal{A}}\Delta_{a}$. The regret analysis depends on several key assumptions stated below.
@@ 5 Regret Analysis
For time $t$ and any intervention $a$, we have $\left\|\phi(a,X_{t})\right\|\leq 1$. In addition, we have $\|\gamma^{*}\|\leq 1$ and $\|\beta^{*}\|\leq 1$.
@@
Assumption 1 is used to make the regret bounds scale-free for convenience; otherwise, we may assume $\left\|\phi(a,X_{t})\right\|\leq C$, $\|\gamma^{*}\|\leq C$, and $\|\beta^{*}\|\leq C$ instead, and the regret bound would increase by a factor of $C$. In practice, this assumption can also be satisfied by scaling the data. Assumption 2 restricts the magnitude of the proximal outcome, which is adapted from the generalized linear bandit literature (Filippi et al., 2010). In Lemma S1.1, S1.3, and S1.4, presented across Sections S3.1, S3.3, and S3.4 of the Supplement, we demonstrate that Assumption 2 validates the sub-Gaussian nature of the noise term in each model. We make an additional Assumption 3 to facilitate the regret analysis for TS-NB and TS-ZINB.
Theorem 5.1.
Suppose we run TS-NB for a total of $T$ time points. Under Assumptions 1-3 (also satisfying certain mild regularity conditions defined in the Supplement), the Bayesian regret of TS-NB is upper-bounded as
$\mathcal{R}_{\text{Bayes}}(T)$$\leq$ $\sqrt{d}+\mathcal{O}(1)+\left[2sY_{\max}\kappa_{\min}^{-1}\sqrt{d\log(T/d)+4\log T}\right]\sqrt{2dT\log\left(\frac{T}{d}\right)},$
@@
Theorem 5.1 establishes $\tilde{O}(d\sqrt{T})$ Bayesian regret bound, which coincides with the Bayesian regret bound for generalized linear TS (Russo and Van Roy, 2014). A detailed proof of Theorem 5.1 is deferred to Section S1.1 of the Supplement.
Theorem 5.2.
Suppose we run TS-ZIP for a total of $T$ time points. Under Assumptions 1-3 (also satisfying certain mild regularity conditions defined in the Supplement), the Bayesian regret of TS-ZIP is upper-bounded as
$\mathcal{R}_{\text{Bayes}}(T)$$\leq$ $\sqrt{d}+\mathcal{O}(1)+\left[2q(Y_{\max}\kappa_{1,\min}^{-1}+\kappa_{2,\min}^{-1})\sqrt{d\log(T/d)+4\log T}\right]\sqrt{2dT\log\left(\frac{T}{d}\right)},$
@@
Theorem 5.2 also establishes $\tilde{O}(d\sqrt{T})$ Bayesian regret bound for TS-ZIP. The proof is deferred to Sections S1.2 of the Supplement. The Bayesian regret bound for TS-ZINB has the same form as TS-ZIP, but under slightly different conditions, and the proof can be found in Section S1.3 of the Supplement.
In the second part of our analysis, we turn our focus to the evaluation of frequentist regret associated with the proposed algorithms. Frequentist regret, in contrast to Bayesian regret, provides a measure of performance under the assumption that true parameters of the model are fixed but unknown, rather than being treated as random variables (Hamidi and Bayati, 2020).
Theorem 5.3.
@@
$\mathcal{R}(T,\beta^{*})$$\leq$ $(\tau+2)\Delta_{\max}+sY_{\max}\kappa_{\min}^{-1}C\left(1+\sqrt{2\kappa_{\min}^{-1}\kappa_{\max}\log(KT)}\right)$
$\times$ $\sqrt{d\log(T/d)+2\log T}\sqrt{2dT\log(T/d)},$
where the number of exploration rounds $\tau$ satisfies $\lambda_{\min}(G_{\tau})\geq 1$. The constant $C$ is in the proof.
Theorem 5.3 establishes $\tilde{O}(d\sqrt{T\log K})$ frequentist regret bound. We defer the proof of Theorem 5.3 to Section S2.1 of the Supplement.
Theorem 5.4.